OpenStudy (anonymous):

The volume of a cylinder is V = pi r2h. Solve for h (height) in this formula.

HanAkoSolo (jamierox4ev3r):

\[V=\pi r^2h\] if you divide this equation by \(r^2h\), you are left with \(h=\frac{V}{\pi r^2h}\)

HanAkoSolo (jamierox4ev3r):

\(\Huge\ h=\frac{V}{\pi r^2} \) perhaps that is a bit clearer :P So you are left with this (on the second one, i accidentallly added an h on the denominator. Please disregard that, thank you)

HanAkoSolo (jamierox4ev3r):

soo @mathematicshead , does that make sense?

OpenStudy (anonymous):

:))))))))))) yessssssssssss

HanAkoSolo (jamierox4ev3r):

okee then :)